# A Principled Framework for Evaluating on Typologically Diverse Languages

### v2

@@ 6.3 Results and Discussion
Figure 6 shows that our methods consistently retrieve more diverse samples than all baselines. For the pairwise metrics (MDP, FVO), this difference is especially large for smaller samples; it is easier to avoid overlap when sampling fewer languages. This is in line with earlier findings from typology, which described the trade-off between coverage and independence (Miestamo, Bakker, and Arppe, 2016). While the FVI difference across methods is small for large sample sizes, we find that our methods retrieve a higher inclusion of feature values for small samples (<20) than the baselines. This is especially relevant for NLP, where multilingual hypotheses are commonly tested on a small set of ‘typologically diverse’ languages (Median = 11; Ploeger et al. 2024). Lastly, we observe that MaxSum and MaxMin consistently retrieve samples with higher feature entropy than other methods.
@@
Next, we compare the methods more in detail, for a specific value of . We zoom in on , since the metrics tend to flatten off for all methods around that value, as seen in Figure 6. We measure the MPD, FVO, FVI and for the samples that all our sampling approaches retrieve. We run the non-deterministic methods 10 times. The results (Table 1) show that both MaxSum and MaxMin sampling retrieve considerably better results than all baselines.
@@ 7 How does Typological Diversity relate to Generalizability?
In this work, we argue that increasing the linguistic representativeness of a language sample leads to better generalizability of results in multilingual NLP. Currently, multilingual language technology is typically only evaluated on a handful of seemingly randomly selected languages. This lack of systematic language sampling makes it difficult to assess how multilingually generalizable such technology really is. In this section, we assess whether sampling typologically diverse languages actually leads to better estimates of general multilingual performance. We apply our sampling framework to three downstream NLP tasks, demonstrating how sampling diverse test languages affects the generalizability of the results. Specifically, we look into natural language understanding (topic classification), natural language generation (machine translation) and subword segmentation.
@@ 7.1 Natural Language Understanding
To assess the extent to which samples drawn with different methods reflect general multilingual performance, it is important we sample from a dataset that includes sufficient typological variation. The Taxi1500 dataset (Ma et al., 2023), which spans more than 1500 languages101010The intersection between Grambank and Taxi1500 contains 1506 languages., covers a large part of the ‘typological space’ (see Appendix B). Importantly, its evaluation set is parallel across languages, which facilitates cross-lingual comparison. The benchmarking task of Taxi1500 is classifying the topic of bible verses, where the six possible topics are recommendation, faith, description, sin, grace and violence. These topics were manually annotated on an English bible, and applied to other languages through annotation projection to create a large-scale parallel benchmark. We compare the results of widely popular classification models (mBERT, Devlin et al., 2019; XLM-R, Conneau et al., 2020) as well as the more recently introduced Glot500-m (Imani et al., 2023).
@@
In Figure 7, we go beyond average results, and inspect the distribution of the results. We compare the distribution of the accuracy over all 1506 languages with samples () from two all methods: MaxSum, Maxmin (ours) and convenience (reflecting common practice in NLP). We observe that the convenience baseline gives consistent overestimations of multilingual performance, while the results of our systematic sampling method consistently reflect the distributions of results for all languages in the dataset. Note that, as illustrated in Figure 1, the objective of systematic language sampling is to draw a language sample from the sampling frame that optimally reflects the larger population. This implies that performance estimates on true typologically diverse languages indeed paint a more generalizable picture of multilingual model performance.
@@ 7.2 Natural Language Generation
Beyond natural language understanding, we evaluate the effect of diverse sampling on a generative task, namely machine translation (MT). We evaluate the effect of the source language on MT performance to English, using the multilingual NLLB-200 model (3.3B parameters).111111huggingface/facebook/nllb-200-3.3B The results are retrieved from the OPUS Dashboard (Tiedemann and de Gibert, 2023). Due to cross-lingual differences, MT performance can only be compared on the same target language and the same test sets. We choose to focus on investigating translation to English specifically, as this is the only language for which the dashboard contains results from all source languages in FloRes-200. English-centric MT has widely been prioritized in MT research, largely motivated by speaker numbers and information access, as argued by Bugliarello et al. (2020), among others. Furthermore, to keep the focus on typological characteristics and avoid the confounding factor of writing systems, we only take into account languages that are written in Latin script. The sampling frame contains 89 source languages in total.
@@
Furthermore, sampling with the MaxSum objective leads to a more conservative performance estimate than the MaxMin objective, consistently highlighting its potential for application in worst-case scenarios.
@@ 7.3 Subword Segmentation
Subword tokenization (Sennrich, Haddow, and Birch, 2016) is an important component in multilingual text processing, which is at the basis of many popular LLMs. Splitting tokens into subwords facilitates better generalizability to languages with complex morphology, and better handles out-of-vocabulary tokens. Yet, tokenizers of popular multilingual models have been shown to retrieve varying results across languages with varying morphosyntactic properties (Gutierrez-Vasques et al., 2021; Gutierrez-Vasques, Bentz, and Samardžić, 2023). Highly synthetic languages such as Finnish pose different tokenization challenges than languages with relatively little morphological complexity such as Dutch. The number of subwords can indicate over-segmentation, which can have far-reaching consequences for a user: for instance, the token-based cost of LLM APIs is higher, and processing more separate tokens introduces latency (Petrov et al., 2023).
@@
Here, we conduct a large-scale cross-lingual sampling comparison of subword tokenization with tokenizers of popular multilingual models. For comparability between languages, the data should ideally be parallel across languages. Previous work (Ahia et al., 2023; Petrov et al., 2023) evaluates on the FLoRes-200 dataset, as this is multi-parallel. We considerably extend upon their language coverage in our analysis, by evaluating on text from the Parallel Bible Corpus (Mayer and Cysouw, 2014), which is the largest massively multi-parallel dataset in terms of language coverage. We first match language ISO codes to Glottocodes. We then select all languages that have Grambank coverage, and control for script (i.e. we filter out non-Latin scripts). This retrieves 571 languages, which is the broadest language coverage thus far in tokenization analysis. This is important, because we aim to test to what extent the sample represent the sampling frame, and potentially the sampling universe. We select the longest bible per language, and sample 1,000 commonly available verses for our evaluation, to retain multi-parallelism. We analyse tokenizers of four popular language models with tokenizers publicly available on HuggingFace121212https://huggingface.co: multilingual BERT (Devlin et al., 2019), LLama 3.3 70B Instruct (Dubey et al., 2024), Qwen2.5-32B (Team, 2024) and Phi 4 (Abdin et al., 2024). Similar to Ahia et al. (2023), we measure the amount of segmentation through the average number of subwords per verse.131313We do not use the fertility measure (Rust et al., 2021), because it assumes similar word tokenizer performance across languages. We do not use ‘premiums’, the disparity of tokenization length between parallel sentences in two languages (Petrov et al., 2023), because these are by definition relative to another language; we instead want to compare across languages directly.
@@
The results are in Figure 9. For all four models, we observe that the convenience baseline retrieves a considerably lower average number of subwords, with a generally smaller range than other methods. This implies that evaluating on the 20 most commonly included languages in ‘typologically diverse’ samples in NLP gives an overly optimistic image of general multilingual tokenizer performance. Yet, MaxMin retrieves averages and spreads that more closely resemble the average and spread of all 500+ languages in the experiment. This suggests that a priori typologically informed language sampling can improve the generalizability of the results, based on a language sample.
@@ 8.1 Guiding Dataset Expansion Efforts
Interestingly, one language stands out as the most diverse expansion language for multiple datasets (FloRes-200, UD and XCOPA): Tariana (language family: Arawakan). From a linguistic typology perspective, this makes sense. Firstly, Tariana is spoken in “a very remote area” in the Vaupés area in the North West of Brasil, which is “not easy to get to” (Aikhenvald, 2003b). Also, Tariana is a polysynthetic language (Aikhenvald, 2003a), a grammatical property that is likely uncommon in popular NLP datasets. Yet, there are more extensive, cultural, reasons for Tariana standing out. The area where the language is spoken, is characterized by “obligatory multilingualism, dictated by the principles of linguistic exogamy” (Aikhenvald, 2003a). This means that marriage only occurs between speakers of different languages. There has been a “strong inhibition against ‘language-mixing’, viewed in terms of lexical loans.”, and the language includes “independent innovations […] divergent from those found in closely related languages.” (Aikhenvald, 2003b). While this case is an indicator that our method indeed achieves high diversity and thus does what we expect, we do not argue that Tariana should then be included in all NLP datasets. Instead, speaker needs and data availability should be taken into account, which can be done by narrowing the sampling frame. We demonstrate such a case in the next section.
@@
Again, this makes sense from a linguistic typology perspective, as Seri is a language isolate which “does indeed have some very special characteristics” (Marlett, 2000). In Table 5, we report the effect of adding Seri to UD v2.14 on all our metrics. We observe that, which the exception of FVI, diversity is improved with respect to all metrics.
@@ 9 Limitations, Ethical Considerations and Discussion
Our framework enables typologically diverse language sampling. Critical assessment of language diversity is vital for the evaluation and development of multilingual language technology that is fair across languages. Still, it should be noted that generalizabilty across typological characteristics constitutes only a fragment of multilingually fair NLP. For instance, Grambank only addresses grammatical phenomena. Such typological databases do not provide cultural information, which may be key for certain research questions (Hershcovich et al., 2022).
Moreover, the feature coverage in typological databases such as Grambank is incomplete, which should be taken into account when drawing conclusions. However, since bibliographic bias in NLP research tends to be much stronger than in typological databases (see convenience baseline, Section 6.3), we believe that information from typological databases can actually enable informed expansion of language coverage.
@@ Acknowledgements.
EP and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. This work received support from the CA21167 COST action UniDive, funded by COST (European Cooperation in Science and Technology). We thank the members of LAGoM NLP group at KU Leuven, the AAU-NLP lab at Aalborg University and the Helsinki-NLP group at the University of Helsinki for feedback on earlier versions of this paper. We thank Robert Östling for pointing us to Anna Sjöberg’s thesis. We thank Kaius Sinnemäki and Anna Sjöberg for discussing the scope of this project with us from a typological perspective. We thank Hedvig Skirgård for helpful feedback on cropping the dataset according to coverage, removing macrolanguages, and the overall scope of this paper. Any remaining errors are our own.References
@@
- Martí et al. (2013) Martí, Rafael, Micael Gallego, Abraham Duarte, and Eduardo G Pardo. 2013. Heuristics and metaheuristics for the maximum diversity problem. Journal of Heuristics, 19:591–615.
@@
- Parreño, Álvarez-Valdés, and Martí (2021) Parreño, Francisco, Ramón Álvarez-Valdés, and Rafael Martí. 2021. Measuring diversity. a review and an empirical analysis. European Journal of Operational Research, 289(2):515–532.

### v3

@@ 6.3 Results and Discussion
Figure 6 shows that our methods consistently retrieve more diverse samples than all baselines. For the pairwise metrics (MDP, FVO), this difference is especially large for smaller samples; it is easier to avoid overlap when sampling fewer languages. This is in line with earlier findings from typology, which described the trade-off between coverage and independence (Miestamo, Bakker, and Arppe, 2016). While the FVI difference across methods is small for large sample sizes, we find that our methods retrieve a higher inclusion of feature values for small samples () than the baselines. This is especially relevant for NLP, where multilingual hypotheses are commonly tested on a small set of ‘typologically diverse’ languages (Median = 11; Ploeger et al. 2024). Lastly, we observe that MaxSum and MaxMin consistently retrieve samples with higher feature entropy than other methods.
@@
Next, we compare the methods in more detail, for a specific value of . We zoom in on , since the metrics tend to flatten off for all methods around that value, as seen in Figure 6. We measure the MPD, FVO, FVI and for the samples that all our sampling approaches retrieve. We run the non-deterministic methods 10 times. The results (Table 1) show that both MaxSum and MaxMin sampling retrieve considerably better results than all baselines.
@@ 7 How does Typological Diversity relate to Generalizability?
In this work, we argue that increasing the linguistic representativeness of a language sample leads to better generalizability of results in multilingual NLP. Currently, multilingual language technology is typically only evaluated on a handful of seemingly randomly selected languages. This lack of systematic language sampling makes it difficult to assess how multilingually generalizable such technology really is. In this section, we assess whether sampling typologically diverse languages actually leads to better estimates of general multilingual performance. We apply our sampling framework to downstream NLP tasks, demonstrating how sampling diverse test languages affects the generalizability of the results. Specifically, we look into natural language understanding (topic classification), natural language generation (machine translation) and subword segmentation. Note that, as illustrated in Figure 1, the objective of systematic language sampling is to draw a language sample from the frame that optimally reflects the larger population. Thus, we evaluate the extent to which the distribution of results in typologically diverse samples reflects the a distribution over a large group of languages.
@@ 7.1 Natural Language Understanding
To assess the extent to which samples drawn with different methods reflect general multilingual performance, it is important to sample from a dataset that includes sufficient typological variation. The Taxi1500 dataset (Ma et al., 2023), which spans more than 1500 languages101010The intersection between Grambank and Taxi1500 contains 1506 languages., covers a large part of the ‘typological space’ (see Appendix B). Importantly, its evaluation set is parallel across languages, which facilitates cross-lingual comparison. The benchmarking task of Taxi1500 is classifying the topic of bible verses, where the six possible topics are recommendation, faith, description, sin, grace and violence. These topics were manually annotated on an English bible, and applied to other languages through annotation projection to create a large-scale parallel benchmark. We compare the results of widely popular classification models (mBERT, Devlin et al., 2019; XLM-R, Conneau et al., 2020) as well as the more recently introduced Glot500-m (Imani et al., 2023).
@@
In Figure 14, we go beyond average results, and inspect the distribution of the results. We compare the distribution of the accuracy over all 1506 languages with samples () from two all methods: MaxSum, Maxmin (ours) and convenience (reflecting common practice in NLP). We observe that the convenience baseline gives consistent overestimations of multilingual performance, while the results of our systematic sampling method consistently reflect the distributions of results for all languages in the dataset. This implies that performance estimates on true typologically diverse languages indeed paint a more generalizable picture of the performance of multilingual models.
@@
Of course, inclusion or exclusion of languages in the training data for these models can be expected to affect the results. In Appendix C, we show the results when sampling only from the pre-training languages of each model. The same pattern holds: the convenience baseline systematically overestimates multilingual performance. Note that in our analysis, we investigated the zero-shot performance of the multilingual models, thus fine-tuning languages are not a confounding factor.
@@ 7.2 Natural Language Generation
Beyond natural language understanding, we evaluate the effect of diverse sampling on a generative task, namely machine translation (MT). We evaluate the effect of the source language on MT performance to English, using the multilingual NLLB-200 model (3.3B parameters).111111huggingface/facebook/nllb-200-3.3B The results are retrieved from the OPUS Dashboard (Tiedemann and de Gibert, 2023). Due to cross-lingual differences, MT performance can only be compared on the same target language and the same test sets. We choose to focus on investigating translation to English specifically, as this is the only language for which the dashboard contains results from all source languages in FloRes-200. English-centric MT has widely been prioritized in MT research, largely motivated by speaker numbers and information access, as argued by Bugliarello et al. (2020), among others. Furthermore, since we do not propose a general language ‘leaderboard’, but rather systematically investigate the influence of typological characteristics on NLP evaluation, we avoid the confounding factor of writing systems. We control for script in our experiments by sampling from languages that use the Latin script specifically, since this maximizes the number of languages in our sampling frame. It should be noted that this limits the generalizability of the findings of this experiment, since we exclude a sizable group of languages.
@@
Furthermore, sampling with the MaxSum objective leads to a more conservative performance estimate than the MaxMin objective Thus, MaxSum sampling is more suitable for application in worst-case evaluation scenarios, whereas the average of MaxMin samples can be expected to better reflect a general distribution.
@@ 7.3 Subword Segmentation
Subword tokenization (Sennrich, Haddow, and Birch, 2016) is an important component in multilingual text processing, which is at the basis of many popular LLMs. While it is not a downstream task, suboptimal tokenization can have large implications for downstream results. Splitting tokens into subwords facilitates better generalizability to languages with complex morphology, and better handles out-of-vocabulary tokens. Yet, tokenizers of popular multilingual models have been shown to retrieve varying results across languages with varying morphosyntactic properties (Gutierrez-Vasques et al., 2021; Gutierrez-Vasques, Bentz, and Samardžić, 2023). Highly synthetic languages such as Finnish pose different tokenization challenges than languages with relatively little morphological complexity such as Dutch. The number of subwords can indicate over-segmentation, which can have far-reaching consequences for a user: for instance, the token-based cost of LLM APIs is higher, and processing more separate tokens introduces latency (Petrov et al., 2023). Previous work (Ahia et al., 2023; Petrov et al., 2023) therefore links segmentation to the concept of multilingual fairness. To extend the scope of our evaluation beyond strict downstream results to fairness more generally, we include a systematic comparison of subword segmenters.
@@
Here, we conduct a large-scale cross-lingual sampling comparison of subword tokenization with tokenizers of popular multilingual models. For comparability between languages, the data should ideally be parallel across languages. Previous work (Ahia et al., 2023; Petrov et al., 2023) evaluates on the FLoRes-200 dataset, as this is multi-parallel. We considerably extend upon their language coverage in our analysis, by evaluating on text from the Parallel Bible Corpus (Mayer and Cysouw, 2014), which is the largest massively multi-parallel dataset in terms of language coverage. We first match language ISO codes to Glottocodes. We then select all languages that have Grambank coverage, and control for script (i.e. we filter out non-Latin scripts). This retrieves 571 languages, which is the broadest language coverage thus far in tokenization analysis. This is important, because we aim to test to what extent the sample represents the sampling frame, and potentially the sampling universe. We select the longest bible per language, and sample 1,000 commonly available verses for our evaluation, to retain multi-parallelism. We analyse tokenizers of four popular language models with tokenizers publicly available on HuggingFace121212https://huggingface.co: multilingual BERT (Devlin et al., 2019), LLama 3.3 70B Instruct (Dubey et al., 2024), Qwen2.5-32B (Team, 2024) and Phi 4 (Abdin et al., 2024). Similar to Ahia et al. (2023), we measure the amount of segmentation through the average number of subwords per verse.131313We do not use the fertility measure (Rust et al., 2021), because it assumes similar word tokenizer performance across languages. We do not use ‘premiums’, the disparity of tokenization length between parallel sentences in two languages (Petrov et al., 2023), because these are by definition relative to another language; we instead want to compare across languages directly.
@@
The results are shown in Figure 9. For all four models, we observe that the convenience baseline retrieves a considerably lower average number of subwords, with a generally smaller range than other methods. This implies that evaluating on the 20 most commonly included languages in ‘typologically diverse’ samples in NLP gives an overly optimistic image of general multilingual tokenizer performance. Yet, MaxMin retrieves averages and spreads that more closely resemble the average and spread of all 500+ languages in the experiment. This suggests that a priori typologically informed language sampling can improve the generalizability of the results, based on a language sample.
@@ 8.1 Guiding Dataset Expansion Efforts
Interestingly, one language stands out as the most diverse expansion language for multiple datasets (FloRes-200, UD and XCOPA): Tariana (language family: Arawakan). From a linguistic typology perspective, this makes sense. Firstly, Tariana is spoken in “a very remote area” in the Vaupés area in the North West of Brasil, which is “not easy to get to” (Aikhenvald, 2003b). Also, Tariana is a polysynthetic language (Aikhenvald, 2003a), a grammatical property that is likely uncommon in popular NLP datasets. Yet, there are more extensive, cultural, reasons for Tariana standing out. The area where the language is spoken is characterized by “obligatory multilingualism, dictated by the principles of linguistic exogamy” (Aikhenvald, 2003a). This means that marriage only occurs between speakers of different languages. There has been a “strong inhibition against ‘language-mixing’, viewed in terms of lexical loans.”, and the language includes “independent innovations […] divergent from those found in closely related languages.” (Aikhenvald, 2003b). While this case is an indicator that our method indeed achieves high diversity and thus does what we expect, we do not argue that Tariana should then be included in all NLP datasets. Instead, speaker needs and data availability should be taken into account, which can be done by narrowing the sampling frame. We demonstrate such a case in the next section.
@@
Again, this makes sense from a linguistic typology perspective, as Seri is a language isolate which “does indeed have some very special characteristics” (Marlett, 2000). In Table 5, we report the effect of adding Seri to UD v2.14 on all our metrics. We observe that, with the exception of FVI, diversity is improved with respect to all metrics.
@@ 9 Limitations, Ethical Considerations and Discussion
Our framework enables typologically diverse language sampling. Critical assessment of language diversity is vital for the evaluation and development of multilingual language technology that is fair across languages. Still, it should be noted that generalizability across typological characteristics constitutes only a fragment of multilingually fair NLP. For instance, Grambank only addresses grammatical phenomena. Such typological databases do not provide cultural information, which may be key for certain research questions (Hershcovich et al., 2022).
Moreover, the feature coverage in typological databases such as Grambank is incomplete, which should be taken into account when drawing conclusions. As such, our evaluation is not completely representative. However, since bibliographic bias in NLP research tends to be much stronger than in typological databases (see convenience baseline, Section 6.3), we believe that information from typological databases can actually enable informed expansion of language coverage. For example, through informed dataset expansion (Section 8.1), we hope our framework can be used to further increase typological representativeness in future datasets.
@@ Acknowledgements.
We thank the members of LAGoM-NLP at KU Leuven, the AAU-NLP lab at Aalborg University, and the Helsinki-NLP group at the University of Helsinki for feedback on earlier versions of this paper. We thank Robert Östling for pointing us to Anna Sjöberg’s thesis. We thank Kaius Sinnemäki and Anna Sjöberg for discussing the scope of this project with us from a typological perspective. We thank Hedvig Skirgård for helpful feedback on cropping the dataset according to coverage, removing macrolanguages, and the overall scope of this paper. Any remaining errors are our own. EP and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. This work received support from the CA21167 COST action UniDive, funded by COST (European Cooperation in Science and Technology).References
@@
- Mart´ı et al. (2013) Martí, Rafael, Micael Gallego, Abraham Duarte, and Eduardo G Pardo. 2013. Heuristics and metaheuristics for the maximum diversity problem. Journal of Heuristics, 19:591–615.
@@
- Parreño, Álvarez-Valdés, and Mart´ı (2021) Parreño, Francisco, Ramón Álvarez-Valdés, and Rafael Martí. 2021. Measuring diversity. a review and an empirical analysis. European Journal of Operational Research, 289(2):515–532.